OpenStudy (anonymous):

Find the vertices and foci of the ellipse. 25x^2 + y^2 = −2y Not sure how to put this is the correct format

Nnesha (nnesha):

\[\huge\rm \frac{( x+h )^2}{a^2 } + \frac{ (y+k)^2 }{ b^2}=1\] ellipse equation so first step) set that equation equal to 1

OpenStudy (anonymous):

So I divide it all by 2y?

Nnesha (nnesha):

nope .... 2y is negative at the right side so add 2y then use completing the square method

OpenStudy (anonymous):

I got the smaller vertices points: (0,-2) and (0,0) But cannot find the foci...

Nnesha (nnesha):

is it horizontal or vertical ??

Nnesha (nnesha):

for vertical \[\huge\rm (h ,k \pm c)\]

Nnesha (nnesha):

a^2 = ? b^2 = ? how did you get 26 at the numerator ?

Nnesha (nnesha):

remember ( h,k) is the center of the ellipse so formula to find foci is \[\huge\rm (h , k \pm c)\] what is the center ?? that you need to find foci